# Role of symmetry in the correlated orientational disorder phase of entropic crystals

### v1

@@ Abstract
Correlated orientational disordered phase is a thermodynamic phase with the constituent particles following some conserved laws or correlation despite the orientational disorder present in the system. Polyhedral particles in the presence of hard interaction serve the purpose of studying the orientational disorder in entropy driven self-assembly. We report the computational evidence of the role of symmetry of both particle and self-assembled crystal in the single-component system of anisotropic polyhedra in the correlated orientational disordered phase. The point groups of the particle and crystal structure appear to make a direct correspondence with the existence of multiple unique orientations of the system satisfying all the orientational conserved quantities towards the correlation. As a signature of correlation, the system exhibits orientationally frozen state at sufficiently higher packing fractions and discrete orientational hopping of the particles are observed at comparatively lower density solids. This symmetry relationship appears to control all the aspects of orientational correlation in the crystalline solids which might be useful to draw fundamental insights about the disordered crystalline solids with orientational correlation as well as designing the disorder in the crystals.
@@
Ordering of classical many-body systems is a spontaneous process in nature which occur in different length scales Whitesides and Boncheva (2002). Crystallization drives a disordered system to a translationally ordered state where all the constituents form a translational broken symmetry state as observed in the domain of colloidal crystals Pawel (1983); Dinsmore et al. (1998); Li et al. (2016); Orr et al. (2022); Boles et al. (2016), molecular crystals Timmermans (1961); Guihrie and Mccullough (1961); Reynolds (1975); Harfenist et al. (1996); Hochli et al. (1990); Nitta (1959) or nanoparticle superlattices Mirkin et al. (1996); Macfarlane et al. (1996); Whitesides and Boncheva (2002); Maye et al. (2007); Henzie et al. (2012); Bodnarchuk et al. (2011); Boneschanscher et al. (2014); Lin et al. (2017); Zhou et al. (2022); Geuchies et al. (2016); Shevchenko et al. (2006); Zhang et al. (2013); Jones et al. (2010); Glotzer and Solomon (2007). Crystal structure of rigid anisotropic particles exhibit different kinds of orientational phases due to the rotational motion of the particles while maintaining proper translational order Mirkin et al. (1996); Glotzer and Solomon (2007); O’Brien et al. (2016); Damasceno et al. (2012a); Gantapara et al. (2013). The possible orientational phases in the crystalline assemblies include the plastic crystals Timmermans (1961); Reynolds (1975); Lee et al. (2023); Karas et al. (2019); Dullens and Petukhov (2007); Meijer et al. (2017), orientationally ordered crystals Damasceno et al. (2012a); Karas et al. (2019); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012), correlated disordered crystals Keen and Goodwin (2015); Chaney et al. (2021); Meekel and Goodwin (2021) or orientational glass Loidl (1989); Hochli et al. (1990); Karas et al. (2019); Abbas et al. (2022); Vdovichenko et al. (2015). The existence of these orientational phases were reported at different length scales either in experiments Abbas et al. (2022); O’Brien et al. (2016); Jones et al. (2010) or computer simulations Damasceno et al. (2012a); Gantapara et al. (2015); John et al. (2008); Karas et al. (2019). Rotator or plastic crystal is a orientational disorder phase with a long-range translational order and very short-range orientational order Timmermans (1961); Reynolds (1975). All the particles can rotate freely in the orientational space while staying at the lattice sites of the crystal; such states can be termed as “freely rotating plastic crystals”. This phenomenon is well documented in the field of colloidal crystals Pawel (1983); Orr et al. (2022); Boles et al. (2016), molecular crystals Vdovichenko et al. (2015); Even et al. (2016); Beake et al. (2017); Hochli et al. (1990) or nanocrystal superlattices Shevchenko et al. (2006); Zhang et al. (2013); Damasceno et al. (2012b); Maye et al. (2007); Lee et al. (2023); Damasceno et al. (2012a). The existence of orientationally ordered crystal structure with the long-range positional and orientational order was also established in earlier literature Damasceno et al. (2012a); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012). Direct experimental observations of orientational order and disorder have been investigated in colloidal crystals Dullens and Petukhov (2007); Meijer et al. (2017) and nanocrystal superlattices Deng et al. (2020); Elbert et al. (2021); Abbas et al. (2022) very recently and studying the orientations of the particles in the crystal structures became one of promising fields in the contemporary research. Correlated orientationally disordered phase is another kind of orientational phase which is less explored compared to the former two phases Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). The existence of the orientational correlation in the disordered crystalline solids was already reported in the molecular crystals Keen and Goodwin (2015) and became an intriguing field to realize the orientational correlation coupled with the translational order.
@@
Correlated orientational disordered phase was realized as a disordered crystal with multiple unique orientations occurring at the comparatively higher packing fractions and maintaining some nontrivial forms of the correlation Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). Multiple unique orientations of the particles were already reported in the colloidal crystals or nanocrystal superlattices in both experiments Abbas et al. (2022) and computer simulations Shen et al. (2019); Karas et al. (2019); Lee et al. (2023) but lacked the insight of correlation. Our previous computational study reported the existence of few orientational attributes which were preserved in the disordered crystalline solids. We argued the collective occurrence of all orientational attributes as the signature of strong orientational correlation Kundu et al. (2023). The phase persisted within certain range of packing fractions, where the orientational motion of the particles were frozen at the densest states and discretely mobile within the unique orientations at comparatively lower packing fractions. The discrete hopping of anisotropic polyhedra within the unique orientations were observed multiple times in the crystalline solids of hard polyhedra in both two Shen et al. (2019) or three dimensions Lee et al. (2023). We also anticipated the “symmetry mismatch” between the non-crystallographic point group of the polyhedral particle and point group of the crystallographic symmetry responsible for the existence of the correlated disordered phase. This gave rise of an obvious question whether the point groups of the particle and crystal could be coupled in some forms to get a better handle of the orientational correlation under the translational broken symmetry state in three dimensions.
@@ IV Conclusion
We report the computational evidence of the role of both particle point group and crystallographic point group in controlling the correlated orientational disorder phase in the entropic crystals. The disorder is maintained in the crystalline system without any spatial arrangements where a few orientational attributes are conserved following some complex forms of correlation. In this research, we discuss the orientational correlation from the symmetry aspect of both the particle and crystal structure. The appearance of the finite number of unique orientations in the system was realized in the context of particle and crystallographic point group. For EPD system, the number of unique orientations appeared to be four where as the axes of the EPD shapes were aligned in the directions of eight crystallographic axes. As the point group of EPD shape contains a plane, the shape is identical about the plane perpendicular to the axes of the shape. So, two axes separated by are considered as orientationally equivalent. As a result, two opposite directions of axes of the EPD shape separated by , are equivalent. Consequently, the number of unique orientations turns down to four from eight for EPD shape. This argument is straightforward and can be verified for the EPG shape containing point group with no plane leading to the eight unique orientations in the system. For the ESG shape with point group not containing any plane, the number of unique orientations appeared to be six, which was also consistent with the alignment of the axes of the particles with six axes of the crystal structure. The possible unique orientations of the shapes in the correlated disordered phase can be described by combining the point groups of both the particle and crystal structure. One obvious interpretation of the equal compartmentalization of the particles within the unique orientations is, the equal probability of the crystallographic symmetry axes in defining the alignments of the particles in the unit cells. As a particular kind of rotational symmetry axes of the crystal structure are separated by fixed angles, the directions of the highest order particle symmetry axes maintain certain angular differences in three dimensions leading to a fixed difference in the quaternions angles appearing in the distribution of all pairwise angles as shown in the third row of Fig. 1. All the conserved attributes corresponding to the correlated disordered phase can be explained in terms of the point group of both particle and crystal structure. For the shapes reported in this study, there exists only two directions of highest order rotational symmetry axes for the particle if the two opposite directions separated by are considered as non-equivalent i.e., axis for EPD and EPG shape and axis for ESG shape. The shapes having multiple highest order particle symmetry axes will be considered in the future investigations. Here, it turns out to be any particular kind of crystallographic axes (any of the , and for the cubic structure) responsible in the alignment of the particles. In general, the alignments will involve multiple kinds of crystallographic axes if large number of unique orientations occur in the system and the possibilities are open for further studies.

### v2

# Role of symmetry in the correlated orientational disordered phase of entropic crystals

@@ Abstract
The crystalline solids with lack of orientational ordering serve the purpose of studying the disordered systems with many fundamental applications in contemporary research. Despite the orientational disorder, the constituent particles sometime follow some conserved laws or correlation while maintaining the translational order depicted by the crystal structure. Such phase is called “correlated orientational disordered” phase in the crystal. Polyhedral particles in the presence of hard-core interaction provide a great opportunity to study the orientational order and disorder in entropy driven self-assembly. We report the computational evidence of the role of symmetry of the polyhedral particles and cubic crystalline structures exhibiting the correlated disordered phases, which are obtained from the simulations of single-component systems. The point groups of the particle and crystal structure appear to be directly connected with the existence of multiple unique orientations satisfying a few other orientational features. In this phase, the systems exhibit orientationally frozen states at sufficiently higher packing fractions and discrete orientational hopping at comparatively lower ones while all the signatures of the correlation are preserved. This symmetry relationship appears to control all the aspects of orientational correlation in the crystalline solids which might be useful to draw fundamental insights about the disordered phases with orientational correlation as well as designing the disorder in the crystals.
@@
Ordering of classical many-body systems is a spontaneous process in nature which occur in different length scales Whitesides and Boncheva (2002). Crystallization drives a disordered system to a translationally ordered state where all the constituents form a translational broken symmetry state as observed in the domain of colloidal crystals Pawel (1983); Dinsmore et al. (1998); Li et al. (2016); Orr et al. (2022); Boles et al. (2016), molecular crystals Timmermans (1961); Guihrie and Mccullough (1961); Reynolds (1975); Harfenist et al. (1996); Hochli et al. (1990); Nitta (1959) or nanoparticle superlattices Mirkin et al. (1996); Macfarlane et al. (1996); Whitesides and Boncheva (2002); Maye et al. (2007); Henzie et al. (2012); Bodnarchuk et al. (2011); Boneschanscher et al. (2014); Lin et al. (2017); Zhou et al. (2022); Geuchies et al. (2016); Shevchenko et al. (2006); Zhang et al. (2013); Jones et al. (2010); Glotzer and Solomon (2007). Crystal structure of anisotropic particles exhibit different kinds of orientational phases due to the rotational motion of the constituent while maintaining proper translational order Mirkin et al. (1996); Glotzer and Solomon (2007); O’Brien et al. (2016); Damasceno et al. (2012a); Gantapara et al. (2013). The possible orientational phases in the crystalline assemblies include the plastic crystals Timmermans (1961); Reynolds (1975); Lee et al. (2023); Karas et al. (2019); Dullens and Petukhov (2007); Meijer et al. (2017), orientationally ordered crystals Damasceno et al. (2012a); Karas et al. (2019); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012), correlated disordered crystals Keen and Goodwin (2015); Chaney et al. (2021); Meekel and Goodwin (2021) or orientational glass Loidl (1989); Hochli et al. (1990); Karas et al. (2019); Abbas et al. (2022); Vdovichenko et al. (2015). The existence of these orientational phases were reported at different length scales either in experiments Abbas et al. (2022); O’Brien et al. (2016); Jones et al. (2010) or computer simulations Damasceno et al. (2012a); Gantapara et al. (2015); John et al. (2008); Karas et al. (2019). Rotator or plastic crystal is a orientational disorder phase with a long-range translational order and very short-range orientational order Timmermans (1961); Reynolds (1975). All the particles can rotate freely in the orientational space while staying at the lattice sites of the crystal; such states can be termed as “freely rotating plastic crystals”. This phenomenon is well documented in the field of colloidal crystals Pawel (1983); Orr et al. (2022); Boles et al. (2016), molecular crystals Vdovichenko et al. (2015); Even et al. (2016); Beake et al. (2017); Hochli et al. (1990) or nanocrystal superlattices Shevchenko et al. (2006); Zhang et al. (2013); Damasceno et al. (2012b); Maye et al. (2007); Lee et al. (2023); Damasceno et al. (2012a). The existence of orientationally ordered crystal structure with the long-range positional and orientational order was also established in earlier literature Damasceno et al. (2012a); Gantapara et al. (2013); Lu et al. (2021); Henzie et al. (2012). Multiple unique orientations of the particles in the orientational glass phase have been thoroughly investigated in both experiments Berret1990; Bounds1982; Loidl (1989); Abbas et al. (2022) and computer simulations Karas et al. (2019) since last few decades. Direct experimental observations of orientational order and disorder were reported in colloidal crystals Dullens and Petukhov (2007); Meijer et al. (2017) and nanocrystal superlattices Deng et al. (2020); Elbert et al. (2021); Abbas et al. (2022) very recently. Understanding the orientations of the particles in the crystal structures became one of promising fields in the contemporary research. Correlated orientationally disordered phase is another orientational phase which is less explored compared to the former three phases Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). The existence of the orientational correlation in the disordered crystalline solids was already reported in the molecular crystals Keen and Goodwin (2015) and turned out to an intriguing field to realize the orientational correlation in the crystalline solids.
@@
Correlated orientational disordered phase was identified in the crystalline solids with multiple unique orientations occurring at the comparatively higher packing fractions where some non-trivial forms of the inter-particle correlation were observed Keen and Goodwin (2015); Simonov and Goodwin (2020); Kundu et al. (2023). Discrete orientations of the particles were already reported in the colloidal crystals or nanocrystal superlattices in both experiments Abbas et al. (2022) and computer simulations Shen et al. (2019); Karas et al. (2019); Lee et al. (2023). Multiple computational studies revealed the discrete orientational hopping of the particles; such orientational phase was characterized as “discrete rotator phase” in the crystals Shen et al. (2019); Lee et al. (2023). The investigators pointed out that the essential differences of this phase and freely rotating plastic crystals could be realized in the context of “discrete” and “continuous” orientational motion of the particles respectively while staying at the lattice sites of the crystals. Our previous computational study independently verified the existence of such orientational behavior of the particles in the disordered crystals and enlightened the long-range orientational correlation in the system. We argued that the “discrete rotator” phase could be thought as a “correlated orientational disordered” phase due to the existence of long-range orientational correlation and disorder in the crystals Kundu et al. (2023), reminiscent to the signatures already observed in the molecular crystals Keen and Goodwin (2015). The existing literature indicated the relationship between the point groups of the particle and local environment Shen et al. (2019); Lee et al. (2023). The commonality of distinct rotational operations in the point groups of the particle and local neighbor was considered to analyze the plastic crystals, discrete rotator phase etc. As our earlier investigation suggests, the orientational disorder occurred at the unit cells of the crystal while preserving all the features of correlation Kundu et al. (2023). So, the existence of this phase could not be characterized without the consideration of crystal structure. This gave rise of an obvious question whether any straightforward relationship between the particles and crystal could be inherited to get a better handle of the orientational correlation in the disordered crystal.
@@ IV Conclusion
We report the computational evidence of a direct connection between the particle point group and crystallographic point group in controlling the correlated orientational disorder phase in the entropic crystals. However, the previous studies indicated the relationship between the point group symmetries of the particles and local neighbor Shen et al. (2019); Lee et al. (2023). The existence of plastic crystals and discrete rotator phase were interpreted in the context of the commonality of the operations in the corresponding point groups. The recent investigation on colloidal clathrates pointed out the essential differences among orientational behavior of guest particles based on the symmetry of polyhedral cages Lee et al. (2023). Though the symmetry of the local neighbor was associated with the polyhedral particle to rationalize the discrete orientations, our observation suggested a more robust affinity explaining the discrete orientations of the particles in the cubic unit cells where the phase existed. The disorder was maintained in the crystalline structure without any particular spatial arrangements, rather a homogeneous distribution was noticed for all the unique orientations. In this research, we discuss the orientational correlation from the symmetry aspect of both the particle and crystal structure. The appearance of the finite number of unique orientations in the system was realized in the setting of particle and crystallographic point group. For EPD system, the number of unique orientations appeared to be four where as the axes of the EPD shapes were aligned in the directions of eight crystallographic axes. As the point group of EPD shape contains a plane, the shape is identical about the plane perpendicular to the axes of the shape. So, two axes separated by are considered as orientationally equivalent. As a result, two opposite directions of axes of the EPD shape separated by , are equivalent. Consequently, the number of unique orientations turns down to four from eight for EPD shape. This argument is straightforward and can be verified for the EPG shape containing point group with no plane leading to the eight unique orientations in the system. For the ESG shape with point group not containing any plane, the number of unique orientations appeared to be six, which was also consistent with the alignment of the axes of the particles with six axes of the crystal structure. The possible unique orientations of the shapes in the correlated disordered phase can be described by combining the point groups of both the particle and crystal structure. One obvious interpretation of the equal compartmentalization of the particles within the unique orientations is, the equal probability of the crystallographic symmetry axes in defining the alignments of the particles in the unit cells. As a particular kind of rotational symmetry axes of the crystal structure are separated by fixed angles, the directions of the highest order particle symmetry axes maintain certain angular differences in three dimensions leading to a fixed difference in the quaternions angles appearing in the distribution of all pairwise angles as shown in the third row of Fig. 1. All the conserved attributes corresponding to the correlated disordered phase can be explained in terms of the point group of both particle and crystal structure. For the shapes reported in this study, there exists only two directions of highest order rotational symmetry axes for the particle if the two opposite directions separated by are considered as non-equivalent i.e., axis for EPD and EPG shape and axis for ESG shape. The shapes having multiple highest order particle symmetry axes will be considered in the future investigations. Here, it turns out to be any particular kind of crystallographic axes (any of the , and for the cubic structure) responsible in the alignment of the particles. In general, the alignments will involve multiple kinds of crystallographic axes if large number of unique orientations occur in the system and the possibilities are open for further studies.